Determine the size of the money multiplier when: the ratio of reserves to total deposits is 10 percent; the ratio of noncheckable deposits to checkable deposits is 30 percent; currency held by the nonbank public is 20 percent of checkable deposits; and the ratio of government deposits to checkable deposits is 10 percent. Also determine the money supply if the monetary base is $10 million
Answer #1

let total checkable deposits be x
Noncheckable deposits = 0.3x
govt deposits = 0.1x
total deposits = 0.3x+0.1x+x = 1.4x
currency = 0.2x
C/D = 0.2x/1.4x = 1/7

Money multiplier = (1+1/7)/(1/10+1/7) = 4.7

Money supply = 4.7*10 = 47million